Which of the following has more atoms, 2.00 g of hydrogen gas or 48.00 g

of titanium?


1
Expert's answer
2021-09-04T02:33:24-0400

1 mol contains 6.022×1023 elementary entities of the substance.

n = m / M

M (H2) = 2 g/mol

M (Ti) = 47.9 g/mol

n (H2) = 2/2 =1 mol

n (Ti) = 48.00 / 47.9 = 1 mol

Since each material is present in equal molar amount, the amount of atoms is also equal.